# Partial identification of principal causal effects under violations of principal ignorability

### v1

@@ 4.2 Partial identification regions when principal ignorability fails
Throughout the rest of this paper, we focus on partial identification regions for parameters of interest without considering prior distributions for these parameters. Partial identification regions are equally supported by the observed data, but the posterior is not necessarily flat in these regions, due to the influence of the prior distribution. Clearly, the prior distribution affects the posterior distribution, and if strong prior knowledge is available, then this should be incorporated. We assume that such knowledge is not available, and that flat priors are placed on unidentified parameters. Using $(\beta_{01},\beta_{10})$ and the identifiable parameters in (3), the PCEs of interest can be written as
$\begin{split}PCE(\boldsymbol{u})&=(\beta_{10}-\frac{\psi_{0}}{\sigma_{s0}}+\frac{\sigma_{s1}}{\sigma_{s0}}\rho\beta_{01})(s_{0}-\phi_{0})+\\ &\qquad(\frac{\psi_{1}}{\sigma_{s1}}-\frac{\sigma_{s0}}{\sigma_{s1}}\rho\beta_{10}-\beta_{01})(s_{1}-\phi_{1})+(\mu^{\prime}_{y1}-\mu^{\prime}_{y0}),\end{split}$ (4)
where $\mu^{\prime}_{yt}=\lambda_{t}+\boldsymbol{\beta}_{t}^{T}(\phi_{0},\phi_{1})$ for $t=0,1$, and $(\beta_{01},\beta_{10})$ are subject to the constraints imposed by (2) and (3). Note that $\mu^{\prime}_{yt}$ is identifiable and corresponds to the mean of potential outcomes, excluding the effects of covariates. Details about their identification are provided in Appendix A. The PCEs depend on the principal strata $\boldsymbol{u}=(s_{0},s_{1})$ and on $(\beta_{01},\beta_{10})$, and therefore we aim to study the partial identification region of $(\beta_{01},\beta_{10})$. For simplicity of the expressions for the partial identification regions of $(\beta_{01},\beta_{10})$, we exclude covariates from the theoretical results regarding partial identification regions. Note that in the appendix we show that to incorporate covariates, we only need to condition additionally on $\boldsymbol{X}$ for all conditional variances or correlations found in the partial identification regions
@@
Suppose that $\text{Var}(Y(1)\mid S(1))\geq\text{Var}(Y(0)\mid S(0))$. Then the partial identification region of $\beta_{01}$ is
$\left[-\sqrt{\frac{\text{Var}(Y(0)\mid S(0))}{(1-\rho^{2})\text{Var}(S(1))}},\sqrt{\frac{\text{Var}(Y(0)\mid S(0))}{(1-\rho^{2})\text{Var}(S(1))}}\right],$
@@
$\begin{split}&\left[-\sqrt{\frac{\text{Var}(Y(1)\mid S(1))}{(1-\rho^{2})\text{Var}(S(0))}},-\sqrt{\frac{\text{Var}(Y(1)\mid S(1))-\text{Var}(Y(0)\mid S(0))}{(1-\rho^{2})\text{Var}(S(0))}}\right]\cup\\ &\qquad\left[\sqrt{\frac{\text{Var}(Y(1)\mid S(1))-\text{Var}(Y(0)\mid S(0))}{(1-\rho^{2})\text{Var}(S(0))}},\sqrt{\frac{\text{Var}(Y(1)\mid S(1))}{(1-\rho^{2})\text{Var}(S(0))}}\right],\end{split}$
@@
$\text{Var}(S(0))\beta_{10}^{2}-\text{Var}(S(1))\beta_{01}^{2}=\frac{\text{Var}(Y(1)\mid S(1))-\text{Var}(Y(0)\mid S(0))}{1-\rho^{2}}.$ (5)
Note that if $\text{Var}(Y(0)\mid S(0))\geq\text{Var}(Y(1)\mid S(1))$, one can simply exchange 0 and 1 in Proposition 1 to obtain the corresponding partial identification region. Also note that all terms in the partial identification region are identifiable since we are assuming for now that $\rho$ is known. Also, an interesting consequence of equation (5) is that if the right-hand side is non-zero, then at least one of $\beta_{01}$ and $\beta_{10}$ must be nonzero, indicating principal ignorability is necessarily violated. Proposition 1 shows that both the sign and magnitude of ($\beta_{01},\beta_{10}$) are not identified, and that these regions can be rather large, which will lead to large partial identification regions for the PCEs of interest. Further, we can see an interplay between the size of the partial identification region for the two parameters. When the partial identification region for $\beta_{01}$ is larger due to a large value of $\text{Var}(Y(0)\mid S(0))$, this leads to a reduction in the width of the region for $\beta_{10}$, and vice-versa. This interplay also effectively guarantees that the partial identification region for at least one of the parameters (if not both) is large, and therefore principal causal effects will have wide partial identification regions as well.
@@ 4.3 Assumptions for identification when principal ignorability is violated
Alternative assumptions beyond the principal ignorability assumption have been proposed in the literature, which we can adopt to obtain point identification. One such modeling assumption explored in recent work [37, 4] in our setting amounts to setting $\beta_{00}=\beta_{10},\beta_{01}=0$. This assumption treats $E[Y(0)\mid\boldsymbol{U},\boldsymbol{X}]$ as a baseline characteristic that is shared between the treatment and control groups [37]. This assumption identifies $\rho$ only up to its sign. However, it is typically reasonable to assume $\rho>0$, which enables full identification of $\rho$. Alternatively, one could make the assumption that both $\lambda_{0}=\lambda_{1}$ and $\beta_{01}=0$, which amounts to assuming there is no direct effect of treatment on the outcome and that the unobserved intermediate does not affect the outcome in the control group. Note that both of these possibilities impose two constraints on the parameter space, which allows us to identify both the outcome model parameters and the unknown correlation $\rho$. If $\rho$ is treated as known, or as a sensitivity parameter to be varied, then only one constraint is needed to obtain identification of the outcome model parameters. It is important to emphasize that while these constraints can be placed to obtain identification, they should only be made if they are deemed plausible in the application to which they are utilized. If one can not make such strong assumptions, then one should continue with partial identification. To this end, we now shift focus to incorporating more plausible assumptions that do not obtain point identification, but can reduce the widths of partial identification regions significantly when such strong identifying assumptions are not plausible.
@@ 5 Weaker alternative assumptions of principal ignorability
Strong assumptions like principal ignorability or the assumptions discussed in Section 4.3 can lead to point identification of principal causal effects, but they are untestable and may be overly restrictive in many settings. Without them, however, partial identification regions such as those seen in Proposition 1 can be overly wide and uninformative even in contexts with simple, parametric models. This is expected to persist or be exacerbated in the context of more flexible outcome models. To allow for violations of principal ignorability and reduce the widths of partial identification regions, we propose two weaker assumptions that can be reasoned about in any particular application. The first assumption addresses the issue of unidentified signs in the effects of the unobserved intermediate.
@@
$\text{sign}\Bigg{(}\dfrac{\partial E(Y\mid T,\boldsymbol{U},\boldsymbol{X})}{\partial S(T)}\Bigg{)}=\text{sign}\Bigg{(}\dfrac{\partial E(Y\mid T,\boldsymbol{U},\boldsymbol{X})}{\partial S_{1-T}}\Bigg{)}.$
@@
Suppose that Assumption 6 holds and $\text{Var}(Y(1)\mid S(1))\geq\text{Var}(Y(0)\mid S(0))$. Assume $\beta_{00}\geq 0$ and $\beta_{11}\geq 0$. The partial identification region for $\beta_{01}$ is
$\Bigg{[}0,\sqrt{\frac{\text{Var}(Y(0)\mid S(0))}{(1-\rho^{2})\text{Var}(S(1))}}\Bigg{]},$
@@
$\Bigg{[}\sqrt{\frac{\text{Var}(Y(1)\mid S(1))-\text{Var}(Y(0)\mid S(0))}{(1-\rho^{2})\text{Var}(S(0))}},\sqrt{\frac{\text{Var}(Y(1)\mid S(1))}{(1-\rho^{2})\text{Var}(S(0))}}\Bigg{]},$
@@
Suppose that Assumption 7 holds, $\text{Var}(Y(1)\mid S(1))\geq\text{Var}(Y(0)\mid S(0))$, $\frac{(\text{Var}(Y(0)\mid S(0)))^{2}}{\text{Var}(Y(0))}\geq\frac{(\text{Var}(Y(1)\mid S(1)))^{2}}{\text{Var}(Y(1))}$, and $\beta_{00}>0$. The partial identification region for $\beta_{01}$ is
$\Bigg{[}-\sqrt{\frac{\text{Var}(Y(0)\mid S(0))\text{Cor}^{2}(Y(0),S(0))}{(1-\rho^{2})\text{Var}(S(1))}},\sqrt{\frac{\text{Var}(Y(0)\mid S(0))\text{Cor}^{2}(Y(0),S(0))}{(1-\rho^{2})\text{Var}(S(1))}}\Bigg{]},$
@@
$\begin{split}&\Bigg{[}-\sqrt{\frac{\text{Var}(Y(1)\mid S(1))-\text{Var}^{2}(Y(0)\mid S(0))/\text{Var}(Y(0))}{(1-\rho^{2})\text{Var}(S(0))}},-\sqrt{\frac{\text{Var}(Y(1)\mid S(1))-\text{Var}(Y(0)\mid S(0))}{(1-\rho^{2})\text{Var}(S(0))}}\Bigg{]}\cup\\ &\qquad\Bigg{[}\sqrt{\frac{\text{Var}(Y(1)\mid S(1))-\text{Var}(Y(0)\mid S(0))}{(1-\rho^{2})\text{Var}(S(0))}},\sqrt{\frac{\text{Var}(Y(1)\mid S(1))-\text{Var}^{2}(Y(0)\mid S(0))/\text{Var}(Y(0))}{(1-\rho^{2})\text{Var}(S(0))}}\Bigg{]},\end{split}$
@@
Note that the result depends on $\text{Var}(Y(1)\mid S(1))\geq\text{Var}(Y(0)\mid S(0))$, ${(\text{Var}(Y(0)\mid S(0)))^{2}}\break/{\text{Var}(Y(0))}\geq{(\text{Var}(Y(1)\mid S(1)))^{2}}/{\text{Var}(Y(1))}$, and $\beta_{00}>0$ all being true. We should emphasize that similar bounds can be derived under any combination of directions for these inequalities, but we focus on one such choice for simplicity. By comparing the results in Propositions 1 and 3, it is clear that Assumption 7 reduces the length of the partial identification region substantially. For $\beta_{01}$ the size of the region is reduced by a factor of $|\text{Cor}(Y(0),S(0))|$, which reflects the strength of the association between the outcome and the observed intermediate within the control $(T=0)$ group. When $|\text{Cor}(Y(0),S(0))|$ is very small, the benefits brought by Assumption 7 are large since it prevents assigning overly large values to the effect of the missing intermediate.
@@
$\dfrac{(\text{Var}[Y\mid S,\boldsymbol{X},T=t])^{2}}{\text{Var}(Y\mid\boldsymbol{X},T=t)}\leq\text{Var}(Y(t)\mid S(t),S(1-t),\boldsymbol{X},T=t),$
@@ 6 Simulations and applications
Here, we examine the theoretical results obtained in the previous sections as well as the practical performance of our weakened identification assumptions for estimating principal causal effects. In the first set of simulation studies, we treat $\rho$ as known, and focus on partial identification of principal causal effects of interest under varying assumptions. Then, we explore the identification of $\rho$ under varying assumptions and sample sizes. Throughout, we do not include covariates $\boldsymbol{X}$, but similar results would be obtained if covariates were additionally adjusted for. Under both scenarios, Gibbs sampling is used, with a single MCMC chain run for a total of 25,000 iterations, where the first 5,000 are burn-in, and a thinning interval of 30 is applied. The details of the Gibbs sampler are included in Appendix B.
@@ 7 Extensions to Bayesian nonparametric modeling
where $\epsilon_{i}\sim N(0,\sigma_{y}^{2})$, $\epsilon\perp\!\!\!\!\perp\boldsymbol{U}_{i}$, and $E[Y_{it}\mid S_{i},\boldsymbol{X_{i}},T_{i}=t]=\mu_{yt}(\boldsymbol{X}_{i},S_{i})$. The subscript $c$ implies that $E[\mu_{yct}(\boldsymbol{X}_{i},\boldsymbol{U}_{i})\mid S_{i},\boldsymbol{X}_{i}]=0$, which means it is centered. Therefore we can view $\mu_{yct}(\boldsymbol{X}_{i},\boldsymbol{U}_{i})$ as a nonparametric extension to $\beta_{t,1-t}S(1-t)$ as it controls the degree of violations to principal ignorability, and if $\mu_{yct}(\boldsymbol{X}_{i},\boldsymbol{U}_{i})=0$, then principal ignorability holds. Our results here are general and hold for any nonparametric Bayesian prior placed on these functions, but either Gaussian processes or Bayesian additive regression tree (BART, [7]) priors would likely be used. The latter of which has been used for estimating principal causal effects [25] and is commonly used more generally within causal inference [15, 13, 27].

### v2

@@ 4.2 Partial identification regions when principal ignorability fails
As mentioned previously, we focus on partial identification regions for parameters of interest without considering prior distributions for these parameters. Partial identification regions are equally supported by the observed data in the sense that they yield the same likelihood, but the posterior is not necessarily flat in these regions, due to the influence of the prior distribution. Clearly, the prior distribution affects the posterior distribution, and if strong prior knowledge is available, then this should be incorporated. We assume that such knowledge is not available and that flat priors are placed on unidentified parameters. Using $(\beta_{01},\beta_{10})$ and the identifiable parameters in (2), the PCEs of interest can be written as
where $\mu^{\prime}_{yt}=\lambda_{t}+\boldsymbol{\beta}_{t}^{T}(\phi_{0},\phi_{1})$ for $t=0,1$, and $(\beta_{01},\beta_{10})$ are subject to the constraints imposed by (1) and (2). Note that $\mu^{\prime}_{yt}$ is identifiable and corresponds to the mean of potential outcomes, excluding the effects of covariates. Details about their identification are provided in Appendix A. The PCEs depend on the principal strata $\boldsymbol{u}=(s_{0},s_{1})$ and on $(\beta_{01},\beta_{10})$, and therefore we aim to study the partial identification region of $(\beta_{01},\beta_{10})$. For simplicity of the expressions for the partial identification regions of $(\beta_{01},\beta_{10})$, we exclude covariates from the theoretical results regarding partial identification regions. Note that in the Appendix we show that to incorporate covariates, we only need to condition additionally on $\boldsymbol{X}$ for all conditional variances or correlations found in the partial identification regions.
@@
Suppose that $\text{Var}\{Y(1)\mid S(1)\}\geq\text{Var}\{Y(0)\mid S(0)\}$. Then the partial identification region of $\beta_{01}$ is
$\left[-\sqrt{\frac{\text{Var}\{Y(0)\mid S(0)\}}{(1-\rho^{2})\text{Var}\{S(1)\}}},\sqrt{\frac{\text{Var}\{Y(0)\mid S(0)\}}{(1-\rho^{2})\text{Var}\{S(1)\}}}\right],$
@@
$\begin{split}&\left[-\sqrt{\frac{\text{Var}\{Y(1)\mid S(1)\}}{(1-\rho^{2})\text{Var}\{S(0)\}}},-\sqrt{\frac{\text{Var}\{Y(1)\mid S(1)\}-\text{Var}\{Y(0)\mid S(0)\}}{(1-\rho^{2})\text{Var}\{S(0)\}}}\right]\cup\\ &\qquad\left[\sqrt{\frac{\text{Var}\{Y(1)\mid S(1)\}-\text{Var}\{Y(0)\mid S(0)\}}{(1-\rho^{2})\text{Var}\{S(0)\}}},\sqrt{\frac{\text{Var}\{Y(1)\mid S(1)\}}{(1-\rho^{2})\text{Var}\{S(0)\}}}\right],\end{split}$
@@
$\text{Var}\{S(0)\}\beta_{10}^{2}-\text{Var}\{S(1)\}\beta_{01}^{2}=\frac{\text{Var}\{Y(1)\mid S(1)\}-\text{Var}\{Y(0)\mid S(0)\}}{1-\rho^{2}}.$ (3)
Note that if $\text{Var}\{Y(0)\mid S(0)\}\geq\text{Var}\{Y(1)\mid S(1)\}$, one can simply exchange 0 and 1 in Proposition 1 to obtain the corresponding partial identification region. Also note that all terms in the partial identification region are identifiable since we are assuming for now that $\rho$ is known. An interesting consequence of equation (3) is that if the right-hand side is non-zero, then at least one of $\beta_{01}$ and $\beta_{10}$ must be nonzero, indicating principal ignorability is necessarily violated. Proposition 1 shows that both the sign and magnitude of ($\beta_{01},\beta_{10}$) are not identified, and that these regions can be rather large, which will lead to large partial identification regions for the PCEs of interest. Further, we can see an interplay between the size of the partial identification region for the two parameters. When the partial identification region for $\beta_{01}$ is larger due to a large value of $\text{Var}(Y(0)\mid S(0))$, this leads to a reduction in the width of the region for $\beta_{10}$, and vice-versa. This interplay also effectively guarantees that the partial identification region for at least one of the parameters (if not both) is large, and therefore principal causal effects will have wide partial identification regions as well. Interestingly, in Appendix C we show that for binary intermediates, the magnitudes of both $\beta_{01}$ and $\beta_{10}$ are in fact identifiable, though their signs are not. This shows that the continuous intermediate setting is inherently more challenging than the binary intermediate setting.
@@ 4.3 Assumptions for identification when principal ignorability is violated
Alternative assumptions beyond the principal ignorability assumption have been proposed in the literature, which we can adopt to obtain point identification. One such modeling assumption explored in recent work [37, 4] in our setting amounts to setting $\beta_{00}=\beta_{10},\beta_{01}=0$. This assumption treats $E\{Y(0)\mid\boldsymbol{U},\boldsymbol{X}\}$ as a baseline characteristic that is shared between the treatment and control groups [37]. This assumption identifies $\rho$ only up to its sign, however, it is typically reasonable to assume $\rho>0$, which enables full identification of $\rho$. Alternatively, one could make the assumption that both $\lambda_{0}=\lambda_{1}$ and $\beta_{01}=0$, which amounts to assuming there is no direct effect of treatment on the outcome and that the unobserved intermediate does not affect the outcome in the control group. Note that both of these possibilities impose two constraints on the parameter space, which allows us to identify both the outcome model parameters and the unknown correlation $\rho$. If $\rho$ is treated as known, or as a sensitivity parameter to be varied, then only one constraint is needed to obtain identification of the outcome model parameters. It is important to emphasize that while these constraints can be placed to obtain identification, they should only be made if they are deemed plausible in the application to which they are utilized. If one can not make such strong assumptions, then one should continue with partial identification. To this end, we now shift focus to incorporating more plausible assumptions that do not obtain point identification, but can reduce the widths of partial identification regions significantly when such strong identifying assumptions are not plausible.
@@ 5 Weaker alternative assumptions of principal ignorability
Strong assumptions like principal ignorability or the assumptions discussed in Section 4.3 can lead to point identification of principal causal effects, but they are untestable and may be overly restrictive in many settings. Without them, however, partial identification regions such as those seen in Proposition 1 can be overly wide and uninformative, even in contexts with simple, parametric models. This is expected to persist or be exacerbated in the context of more flexible outcome models. To allow for violations of principal ignorability and reduce the widths of partial identification regions, we propose two weaker assumptions that can be reasoned about in any particular application. The first assumption addresses the issue of unidentified signs in the effects of the unobserved intermediate.
@@
$\text{sign}\Bigg{\{}\dfrac{\partial E(Y\mid T,\boldsymbol{U},\boldsymbol{X})}{\partial S(T)}\Bigg{\}}=\text{sign}\Bigg{\{}\dfrac{\partial E(Y\mid T,\boldsymbol{U},\boldsymbol{X})}{\partial S(1-T)}\Bigg{\}}.$
@@
Suppose that Assumption 6 holds and $\text{Var}\{Y(1)\mid S(1)\}\geq\text{Var}\{Y(0)\mid S(0)\}$. Assume $\beta_{00}\geq 0$ and $\beta_{11}\geq 0$. The partial identification region for $\beta_{01}$ is
$\Bigg{[}0,\sqrt{\frac{\text{Var}\{Y(0)\mid S(0)\}}{(1-\rho^{2})\text{Var}\{S(1)\}}}\Bigg{]},$
@@
$\Bigg{[}\sqrt{\frac{\text{Var}\{Y(1)\mid S(1)\}-\text{Var}\{Y(0)\mid S(0)\}}{(1-\rho^{2})\text{Var}\{S(0)\}}},\sqrt{\frac{\text{Var}\{Y(1)\mid S(1)\}}{(1-\rho^{2})\text{Var}\{S(0)\}}}\Bigg{]},$
@@
Suppose that Assumption 7 holds, $\text{Var}\{Y(1)\mid S(1)\}\geq\text{Var}\{Y(0)\mid S(0)\}$, $\frac{(\text{Var}\{Y(0)\mid S(0)\})^{2}}{\text{Var}(Y(0))}\geq\frac{(\text{Var}\{Y(1)\mid S(1)\})^{2}}{\text{Var}(Y(1))}$, and $\beta_{00}>0$. The partial identification region for $\beta_{01}$ is
$\Bigg{[}-\sqrt{\frac{\text{Var}\{Y(0)\mid S(0)\}\text{Cor}^{2}(Y(0),S(0))}{(1-\rho^{2})\text{Var}\{S(1)\}}},\sqrt{\frac{\text{Var}\{Y(0)\mid S(0)\}\text{Cor}^{2}(Y(0),S(0))}{(1-\rho^{2})\text{Var}\{S(1)\}}}\Bigg{]},$
@@
$\begin{split}&\Bigg{[}-\sqrt{\frac{\text{Var}\{Y(1)\mid S(1)\}-\text{Var}\{Y(0)\mid S(0)\}/\text{Var}(Y(0))}{(1-\rho^{2})\text{Var}\{S(0)\}}},-\sqrt{\frac{\text{Var}\{Y(1)\mid S(1)\}-\text{Var}\{Y(0)\mid S(0)\}}{(1-\rho^{2})\text{Var}\{S(0)\}}}\Bigg{]}\cup\\ &\qquad\Bigg{[}\sqrt{\frac{\text{Var}\{Y(1)\mid S(1)\}-\text{Var}\{Y(0)\mid S(0)\}}{(1-\rho^{2})\text{Var}\{S(0)\}}},\sqrt{\frac{\text{Var}\{Y(1)\mid S(1)\}-\text{Var}\{Y(0)\mid S(0)\}/\text{Var}(Y(0))}{(1-\rho^{2})\text{Var}\{S(0)\}}}\Bigg{]},\end{split}$
@@
Note that the result depends on $\text{Var}\{Y(1)\mid S(1)\}\geq\text{Var}\{Y(0)\mid S(0)\}$, ${(\text{Var}\{Y(0)\mid S(0)\})^{2}}\break/{\text{Var}(Y(0))}\geq{(\text{Var}\{Y(1)\mid S(1)\})^{2}}/{\text{Var}(Y(1))}$, and $\beta_{00}>0$ all being true. We should emphasize that similar bounds can be derived under any combination of directions for these inequalities, but we focus on one such choice for simplicity. By comparing the results in Propositions 1 and 3, it is clear that Assumption 7 reduces the length of the partial identification region substantially. For $\beta_{01}$ the size of the region is reduced by a factor of $|\text{Cor}(Y(0),S(0))|$, which reflects the strength of the association between the outcome and the observed intermediate within the control $(T=0)$ group. When $|\text{Cor}(Y(0),S(0))|$ is very small, the benefits brought by Assumption 7 are large since it prevents assigning overly large values to the effect of the missing intermediate.
@@
$\dfrac{[\text{Var}\{Y\mid S,\boldsymbol{X},T=t\}]^{2}}{\text{Var}(Y\mid\boldsymbol{X},T=t)}\leq\text{Var}\{Y(t)\mid S(t),S(1-t),\boldsymbol{X},T=t\},$
@@ 6 Simulations and applications
Here, we examine the theoretical results obtained in the previous sections as well as the practical performance of our weakened identification assumptions for estimating principal causal effects. In the first set of simulation studies, we treat $\rho$ as known, and focus on partial identification of principal causal effects of interest under varying assumptions. Then, we explore the identification of $\rho$ under varying assumptions and sample sizes. Throughout, we do not include covariates $\boldsymbol{X}$, but similar results would be obtained if covariates were additionally adjusted for. Under both scenarios, Gibbs sampling is used, with a single MCMC chain run for a total of 25,000 iterations, where the first 5,000 are burn-in, and a thinning interval of 30 is applied. The details of the Gibbs sampler are included in Appendix A.
@@ 7 Extensions to Bayesian nonparametric modeling
where $\epsilon_{i}\sim N(0,\sigma_{y}^{2})$, $\epsilon\perp\!\!\!\!\perp\boldsymbol{U}_{i}$, and $E(Y_{it}\mid S_{i},\boldsymbol{X_{i}},T_{i}=t)=\mu_{yt}(\boldsymbol{X}_{i},S_{i})$. The subscript $c$ implies that $E\{\mu_{yct}(\boldsymbol{X}_{i},\boldsymbol{U}_{i})\mid S_{i},\boldsymbol{X}_{i}\}=0$, which means it is centered. Therefore we can view $\mu_{yct}(\boldsymbol{X}_{i},\boldsymbol{U}_{i})$ as a nonparametric extension of $\beta_{t,1-t}S(1-t)$ as it controls the degree of violations of principal ignorability, and if $\mu_{yct}(\boldsymbol{X}_{i},\boldsymbol{U}_{i})=0$, then principal ignorability holds. Our results here are general and hold for any nonparametric Bayesian prior placed on these functions, but either Gaussian processes or Bayesian additive regression tree (BART, [7]) priors would likely be used. The latter of which has been used for estimating principal causal effects [25] and is commonly used more generally within causal inference [15, 13, 27].